OpenStudy (anonymous):

Three vertixes of a convex quadrilateral are fixed. Find the locus of the fourth vertex if the area of the quadrilateral is constant.

OpenStudy (anonymous):

I came across a similar question somewhere. Wouldn't it be a line?

OpenStudy (anonymous):

um theres several answers i think

OpenStudy (anonymous):

what year are u in!

OpenStudy (anonymous):

|1 1| |0 1| |x y| |1 0| |1 1| 0.5(1 + 1 - (x + y)) = ? 0.5(2 - x - y) = ? 1 - 0.5x - 0.5y = ? -0.5y = 0.5x - 1 y = -x + 2 Not completely sure if this is correct, but that's what I got.

OpenStudy (anonymous):

It would be a line most definitely.

OpenStudy (anonymous):

What I did was use an example to see what I got.

OpenStudy (anonymous):

Is this multiple choice?

OpenStudy (anonymous):

nope.. well the answer says theres three cases to consider. first case answer is y=-2k/a k is constant

OpenStudy (anonymous):

I'm sorry...you're gonna have to ask someone else for this one...the only thing I see is 'locus of points' which would mean a type of graph.

OpenStudy (anonymous):